Halliford Corporation expects to have earnings this coming year of $3.092 per share. Halliford plans to retain all of its earnings for the next two years.​ Then, for the subsequent two​ years, the firm will retain 53% of its earnings. It will retain 20% of its earnings from that point onward. Each​ year, retained earnings will be invested in new projects with an expected return of 23.7% per year. Any earnings that are not retained will be paid out as dividends. Assume​ Halliford's share count remains constant and all earnings growth comes from the investment of retained earnings. If​ Halliford's equity cost of capital is 10.6%​,what price would you estimate for Halliford​ stock?
Answer #1

Halliford’s dividend forecast (g = retention rate × return on new investment)

Year

0

1

2

3

4

5

6

Earnings

1).EPS Growth Rate (vs. prior yr)

23.70%

23.70%

12.561%

12.561%

4.74%

2). EPS

$3.092

$3.8248

$4.7313

$5.3256

$5.9945

$6.2787

Dividends

3). Retention Ratio

100%

100%

53%

53%

20%

20%

4). Dividend Payout Ratio

0%

0%

47%

47%

80%

80%

Dividends(2 x 4) $0 $0    $2.2237 $2.5030 $4.7956 $5.0230

From year 5 on, dividends grow at constant rate of 4.74%. Therefore,

P(5) = D6 / (r - gC) = $5.0230 / (10.6% - 4.74%) = $5.0230 / 0.0586 = $85.72

P(0) = [D3 / (1 + r)3] + [D4 / (1 + r)4] + [(D5 + P5) / (1 + r)5]

= [$2.2237 / 1.1063] + [$2.5030 / 1.1064] + [($4.7956 + $85.72) / 1.1065]

= $1.64 + $1.67 + $54.69 = $58.01